

********************************************************************
UARS WEEKLY STATUS REPORT - 12 OCT 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet the SOLSTICE science
objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry.
We are testing the PACOR-A level 0 SUSIM science files.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
HALOE instrument engineering parameters are within nominal operating
range. Science channel outputs are as expected. There has been some delay
in receiving 24 hour files with the new PACOR A system that has delayed
data evaluation and assessment of the HALOE data. The HALOE FOT has been
working with Rich Beck to insure the capture of HALOE events with the
reduced TDRS support. Rich is maximizing TDRS coverage during HALOE events
and supporting the remainder with the tape recorder. However, there is a
concern as to the quality of the captured data, on some days, with the new
PACOR A system that is being addressed.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans
continues with no problems. The Command Mangement
System processing requires Daily Activity Plans each day
for all operational instruments to maintain continuity
and history of the uplinked command procedures.
The MPG distributed the Instrument Schedule to provide test
data for the PACOR-A system from 17 October to 26 November
2001.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
This report covers 29 September 2001 (Orbit 54946, GMT Day 272) through 05
October 2001 (Orbit 55050, GMT Day 278).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 274/0027 54976 Power ON
SUSIM 274/1627 54986 Power ON (P-1)
SUSIM 274/1724 54987 Power ON (P-2)
HALOE 275/0847 54996 Power OFF (TBAT Y-HI)
SUSIM 275/1223 54999 Power ON (P-3)
HALOE 275/1627 55001 Power ON
The beta angle ranged between the angles of 54.1 degrees to 47.6 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 0 degrees on 17 October 2001 (DOY 290).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 to 1.3 degrees C. Additional battery performance data for each
battery on 26 September 2001 and 3 October 2001 are provided in Appendix D.
The clock error ranged from Š11.4 to 11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
„ OUT-OF-LIMITS = 02
„ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
„ FOT = 20
„ HALOE = 04
„ SUSIM = 03
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
„ FOT = 01
Zero (0) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 01/267 - 2287.496874
· Transponder B: 01/263 - 2287.498962
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/277 show the observatory data loss to be 303 hours, 18
minutes, and 28.016 seconds. This shows a data loss of 5 minutes and 32.992
seconds since the last report period. This is a 0.39 percent data loss, which
equals a 99.61 percent data capture for the mission.
The increase of 5 minutes and 32.992 seconds occurred over one (4) days:
· 10.144 seconds on Day 271 due to a Data Hit
· 09.312 seconds on Day 272 due to a Data Hit
· 10.016 seconds on Day 273 due to a Data Hit
· 5 minutes and 03.520 seconds on Day 274 due to No Data Overlap
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has begun the process of implementing RaytheonÕs Eclipse ground system
and continues to work on the TAC replacement system. These new systems will
provide enhanced capability and make up for the loss of some UARS facilities,
which have been shut down due to budgetary reasons. This will be a huge effort
and must be completed as soon as possible in order to provide a low cost method
for flying UARS. The FOT will need assistance from all entities involved to
insure the success of this operation. The estimated time for completion is
April 2002.
The FOT has also begun the determination of the overall requirements necessary
to meet the needs of the new ground system as a whole. The requirements will
be examined to determine which entity can better meet the need. A weekly
meeting has been set up on Wednesday afternoons to discuss these requirements
as well as address any issues involving the implementation of the new ground
system.
MEETINGS
The UARS FOT attended a UARS Monthly Status Meeting on Wednesday, October 03,
2001.
The UARS FOT attended a UARS Operational Plan Meeting on Wednesday, October 03,
2001.
The UARS FOT attended UARS Technical Team Meeting on Thursday, October 04,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 October 2001 (DOY 290). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 October 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
54952 272/1028 01-598 OBC Clk Adj fr 20.6 to 0.0 msec/day
54976 274/0027 01-599 NBTR NOOPS D-274
54976 274/0027 01-600 HALOE Power ON
54978 274/0241 01-601 HALOE AZIMUTH ADJ.
54984 274/1222 01-599 NBTR NOOPS D-274
54984 274/1222 01-602 OBC/CLK Clk Adj 0.0 to 20.6 msec/day
54986 274/1627 01-604 SUSIM Power ON (P-1)
54987 274/1724 01-604 SUSIM Power ON (P-2)
54987 274/1807 01-603 NBTR NOOPS D-275
54995 275/0607 01-603 NBTR NOOPS D-275
54995 275/0704 01-603 NBTR PBK
54996 275/0847 01-605 HALOE Power OFF (TBAT Y-HI)
54999 275/1223 01-604 SUSIM Power ON (P-3)
55001 275/1627 01-606 HALOE Power ON
55002 275/1724 01-603 NBTR NOOPS D-275
55002 275/1807 01-607 NBTR NOOPS D-276
55011 276/0746 01-607 NBTR NOOPS D-276
55011 276/0848 01-607 NBTR NBTR PBK
55013 276/1110 01-608 OBC/CLK Clk Adj 20.6 to 0.0 msec/day
55015 276/1443 01-607 NBTR NBTR PBK & RTS-58
55018 276/1904 01-609 NBTR NOOPS D-277
55025 277/0607 01-609 NBTR NOOPS D-277
55025 277/0703 01-609 NBTR PBK
55028 277/1111 01-609 NBTR NOOPS D-277
55032 277/1807 01-610 NBTR NOOPS D-278
55039 278/0524 01-610 NBTR NOOPS D278
55047 278/1806 01-612 NBTR NOOPS D279
_________________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 October 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
HATBAT Y-HI Telescope Barrel Temp 54995-6, 55036-40
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 October 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
54952 272/1028 Clk Adj fr 20.6 to 0.0 msec/day
54984 274/1238 Clk Adj 0.0 to 20.6 msec/day
55013 276/1115 Clk Adj 20.6 to 0.0 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 03 October 2001
26 September 2001 - GMT Day 269 Beta = 53.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.36/-3.99 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-22.4 +3.23/+2.60 +13.6/-13.2 28.3 1.020 17.8
3 +11.2/-22.4 +2.60/+1.66 +13.6/-12.8 28.3 1.010 16.7
03 October 2001 - GMT Day 276 Beta = 50.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/-50.0 11.7 0.000 0.0
2 +00.0/-28.0 +4.51/+3.23 +25.2/-50.0 26.1 1.040 19.5
3 +22.4/-22.4 +3.55/+2.29 +24.0/-50.0 26.1 1.020 19.2
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 October 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-059 FOT 55050 278/2310 NA Telemetry Hits
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-054 FOT see list Sept. 01 14170 Generic Late Acquisition
01-059 FOT 55050 278/2310 NA Telemetry Hits
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
SEVNTFW01/40